• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 14:53:42 on March 20, 2012

    Entry number 367192

    keyword=g2p/gep Tuesday SWING, Wed OWL, Wed DAY cancelled; Wed SWING+ on standby (Brief status update)

    g2p/gep shifts on Tuesday SWING, Wed OWL and Wed DAY are canceled. Wed SWING shift and future shifts are on standby.

    This is actually good news because it is a side effect of promising repair efforts on the RHRS septum. The septum will not be able to run at full field, but it looks likely that the bad coils can be bypassed. As this is a relatively small subset of the total windings, the impact on the septum field (and associated reduction in the RHRS maximum momentum acceptance) should be manageable and allow productive use of the RHRS going forward.

    Repair work and associated performance checks on the septum hardware will continue today through day shift Wednesday. Those on shift Wednesday SWING and later should watch the Halog and their email for an update on shift status tomorrow.

    A copy of this log entry has been emailed to: g2p@jlab.org